OpenStudy (babynini):

Exponential equation (log): 625^x=1/√5 Please explain the process for how to solve this!

OpenStudy (asnaseer):

First note that \(625=5^4\). Then also notice that \(\frac{1}{\sqrt{5}}=\frac{1}{5^{0.5}}=5^{-0.5}\) So your equation reduces to:\[(5^4)^x=5^{-0.5}\]Which you should be able to solve without using logs.

OpenStudy (babynini):

You're amazing, thanks so much. Only..how do I solve from there? :)

OpenStudy (babynini):

Nevermind! I got it now.